a garden 7 m by 12 m will be expanded by planting a border of flowers. the border will be of the same width around the entire garden and has an area of 92 m^2
Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52781)

.
No question ----------------------------> no answer.

        No question ----------------------------> no answer.

                No question ----------------------------> no answer.



It is THE RULE of this forum.



In addition, it is the extreme degree of impoliteness to the tutors to submit the posts without questions.



There are more definitive words for it, but they are out of "politecorrectness" frames, so I do not want to use them.


Keep it in your mind when you post to this forum.


Thanks.



Answer by Boreal(15235)

Draw this out and the area of the entire garden+border is (7+2x)(12+2x)=92
84+38x+4x^2=92
4x^2+38x-8=0
2x^2+19x-4=0
x=(1/4)(-19+/-sqrt (361+32)), sqrt (393)=19.82
only positive root is x=0.82/4 meter or 0.205 m or 20.5 cm width
7.41*12.41=91.96 m^2, rounding error.